OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x,y) 8x-y=31 5x+3y=23

OpenStudy (binary3i):

8x-y=31 y=8x-31put this in second equation 5x+3y=23 5x+3(8x-31)=23 29x= 23+93= 116 x=116/23

OpenStudy (anonymous):

So it would be (116,23)? Or (23,116)?

OpenStudy (binary3i):

c'mon its just x now put x in any of the equation and get your y.

OpenStudy (anonymous):

sfr x in top plug it into x in bottom eq take the result and plug it back into top

OpenStudy (anonymous):

Hi, Kitty. Here's the process for doing problems like this. Pick your favorite equation and your favorite variable in that equation. Solve for that variable. Substitute into the OTHER equation for that variable, which will give you an equation that only has one variable in it. Solve for that variable, and you're halfway done! Now that you know one variable, you can look back at either of your first two equations, substitute that variable in and find out what the other variable is. =D